OpenStudy (anonymous):

If r = 12.5 and s = |−13|, which of the following inequalities is not true? r ≠ s r < s r ≥ s r ≤ s

OpenStudy (solomonzelman):

do you know what it means to say that: \(\large\color{black}{ r \ne s }\) ?

OpenStudy (solomonzelman):

can you exclude any of the options?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

Okay, symbols: " \(\large\color{black}{\ne }\) " means that it is "not equal to". For example when we say: \(\large\color{black}{ 3 \ne 4 }\) all we say is that "3 is not equal to 4".

OpenStudy (solomonzelman):

Do you know what the absolute value is? (written as \(\large\color{black}{ \left|~~~~\right| }\) )

OpenStudy (solomonzelman):

in your case it is \(\large\color{black}{ \left|-13~\right| }\) , do you know what this means?

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

thanks? I thought you aren't done yet.

OpenStudy (anonymous):

it means -13?

OpenStudy (solomonzelman):

not exactly.

OpenStudy (anonymous):

or is it 13?

OpenStudy (anonymous):

don't just guess

OpenStudy (solomonzelman):

Yes. When you say an absolute value of the number, you are taking a positive value of the number like (considering positive number a, where "-a" is thus a negative number) \(\large\color{black}{ \left|~-a~\right| =a }\) For example: \(\large\color{black}{ \left|~-5~\right| =5 }\) and \(\large\color{black}{ \left|~-10~\right| =10 }\) the positive inside the absolute value though, remains positive: For example. \(\large\color{black}{ \left|~7~\right| =7 }\)

OpenStudy (anonymous):

so it takes what ever is inside and makes it positive!

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So, if it is : \(\large\color{black}{ \left|~-13~\right| }\) (you can use my code, `\(\large\color{black}{ \left|~-13~\right| }\)` copy paste from the gray) what will it be equal to?

OpenStudy (anonymous):

thanks, I do all my work on the computer and I have no teachers just a superviser

OpenStudy (solomonzelman):

okay, sorry for insisting, but what will \(\large\color{black}{ \left|~-13~\right| }\) be equal to?

OpenStudy (anonymous):

13!

OpenStudy (solomonzelman):

yes, but don't put an exclamation mark, because it means a factorial. I know you just mean to say 13 though, not a factorial.

OpenStudy (solomonzelman):

Anyways: you had: \(\large\color{black}{ s=\left|~-13~\right| }\) and \(\large\color{black}{ r=12.5 }\) now you have: \(\large\color{black}{ s=13 }\) and \(\large\color{black}{ r=12.5 }\)

OpenStudy (anonymous):

sorry im exited, I have a learning disability in math so when I get something it makes me exited

OpenStudy (anonymous):

what does ≥ mean? how different is it from <?

OpenStudy (solomonzelman):

it is fine, and you aren't disabled. but if you keep thinking that why you might imbue that to yourself... anyway, the problem NOW is: If, \(\large\color{blue}{ s=13 }\) and \(\large\color{blue}{ r=12.5 }\) which of the following inequalities is not true? \(\large\color{black}{ r \ne s }\) \(\large\color{black}{ r < s }\) \(\large\color{black}{ r \ge s }\) \(\large\color{black}{ r \le s }\)

OpenStudy (solomonzelman):

you mean how \(\large\color{black}{ \ge }\) is different from \(\large\color{black}{ > }\) ?

OpenStudy (anonymous):

\[r \le s\] right?

OpenStudy (solomonzelman):

you know what this statement means?

OpenStudy (anonymous):

r is less than or equal to s

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

but it is not the answer, because: \(\large\color{black}{ 12.5\le 13}\) is not a true statement.

OpenStudy (solomonzelman):

ohh yes right!

OpenStudy (solomonzelman):

I thought it was asking for a true statement, lol

OpenStudy (anonymous):

thanks for the help but there is some one I want to help so I will have questions l8r if u want to help.